OpenStudy (anonymous):

Jar B has 20% more marbles than jar A. What percent of marbles should be moved from jar A to jar B so that jar B will double jar A in number of marbles

OpenStudy (anonymous):

let x be the percent of marbles removed from jar A, then B = 1.2A B + Ax = 2(A - Ax ) 1.2A + Ax = 2(A - Ax) 1.2A + Ax = 2A - 2Ax Ax + 2Ax = 2A - 1.2A x(3A) = 0.8A x = 0.8A/3A x = 0.26666.... so about 26.67%